# PTM-Psi on the Cloud

### v1

@@ 15                CYX19, CYX29 SSG
Table 1: Each reference number corresponds to a specific case with a particular set of post-translational modifications (PTMs) in specific chain locations. “CYS” denotes a free cysteine residue, while “CYX” represents part of a disulfide bridge in PDB ID 6GVE (i.e. Ref 0) – there are two disulfide bridges in PRK and in CP12, respectively. Abbreviations of PTMs on a specific cysteine include SNO (S-nitrosylation), SSH (S-sulfhydration), SSG or (S-glutathionylation).
@@ 2.3.1 Univariate Metrics
RMSD quantifies the deviation of instantaneous atomic positions from a reference (average) structure:
@@
where $\mathbf{r}_{k}^{\rm ref}$ is the average position of atom $k$ over all frames.
@@
RMSF describes per-residue flexibility. For residue $i$, using backbone heavy-atom center-of-mass $\mathbf{R}_{i}(t)$,
@@
Secondary Structure (DSSP)
Secondary structure assignment tracks helix, sheet, and coil fractions over time using the Dictionary of Secondary Structure in Proteins (DSSP) algorithm23, 24.
@@ 2.3.3 Multivariate Analysis
Multivariate analysis (MVA) refers to the statistical analysis of data that involves multiple variables or “features”. This type of analysis is essential when dealing with complex and massive datasets where each observation consists of multiple measurements. We deployed MVA is to understand the relationships between the variables in the datasets. We used the technique of principal component analysis (PCA) to reduce the dimensionality of the datasets by transforming them into a new set of orthogonal variables that capture the most variance. By pooling measurements per frame into one feature matrix, we revealed correlated motions invisible to one-dimensional plots by applying principal-component analysis (PCA) that reduces the redundancy among overlapping metrics, and transformed the simulations from all instances under various PTM perturbations into a common coordinate system25. By ranking the variability, we can rank the impact of PTM perturbations on the conformational ensemble in the simulations.
Feature extraction: For each reference $i$ in Table 1, there are $N\approx 1000$ samples. At timeframe $t$, we compute an $M$-dimensional feature vector from a selection of atoms that is either from the whole protein complex or from the reduced representation as described in the previous subsection:
@@
Here, $R_{g}$ is the radius of gyration (in nm). $\mathrm{RMSD}$ is the root-mean-square deviation (in nm). $B$ is the average B-factor (in nm2), calculated by
@@
where $\langle\mathrm{RMSF}^{2}\rangle$ denotes the average squared root-mean-square fluctuation per residue in a selected representation. $\phi$ and $\psi$ are the average backbone dihedral angles (in radians). $\mathrm{SS}_{\mathrm{helix}}$, $\mathrm{SS}_{\mathrm{sheet}}$, and $\mathrm{SS}_{\mathrm{coil}}$ are the fractional secondary-structure propensity. $\mathrm{n}_{\mathrm{PTM\text{-}Residue}}$ denotes the count of residual contacts between the PTMed residue and its neighboring residues within 3.0 Å.
@@
and concatenating all of 16 references (e.g., Ref 0 and 15 other PTM perturbations in Table 1):
@@
and we retain the $K$ number of principal components such that the cumulative variance is preserved as follows,
@@
Ranking PTM perturbation: the mean projection onto the first principal component (PC1) for reference $i$ is given by
[…]
@@
measures the PTM perturbations on the conformational ensemble from Ref #i along the dominant mode of variation relative to those of Ref 0. The magnitude $|\Delta\langle\mathrm{PC1}\rangle_{i}|$ indicates deviation from the unmodified reference ensemble, and the sign reflects the direction of the shift in collective conformational space. References are ranked by $\Delta\langle\mathrm{PC1}\rangle_{i}$.
[…]
Here, $v_{1}=[\ell_{1},\ell_{2},\dots,\ell_{M}]^{\mathsf{T}}$ and each loading $\ell_{k}$ informs the strength of the feature $k$ driving PC1 in either a positive “+” or a negative “–” direction.
2.3.4 Kinetic Transition Network (KTN)
The KTN formalism transforms a high-dimensional, noisy molecular trajectory into a simplified yet informative network model that captures both the thermodynamics (through stationary distributions) and kinetics (through transition rates and mean first passage time, MFPTs) of the system. We employed a Kinetic Transition Network (KTN), a graph-based model26, to estimate transition probabilities between metastable states in a conformational ensemble and to identify representative states for each basin. It captures the dynamic connectivity for visualizing and quantifying the kinetics of conformational switching between basins. For this study, we reduced the dimensionality of its feature matrix $X\in\mathbb{R}^{N\times M}$ from Equation (6) by projecting each frame onto the top two principal components. This step transforms the high-dimensional feature space into an interpretable 2D subspace that preserves the most dominant collective motions:
@@ 2.3.4 Kinetic Transition Network (KTN)
where $\mu_{j}\in\mathbb{R}^{2}$ is the centroid of cluster $j$, and $k$ is a user-defined parameter for defining the number of clusters (typically $k=5$) representing the number of coarse-grained states. Each frame $t$ is assigned a discrete cluster label $m_{t}\in\{0,1,\dots,k-1\}$, identifying the kinetic basin it belongs to.
@@
In our KTN figures, the node area scales with $\pi_{j}$, and the label “$p$” on each node is the stationary probability $\pi_{j}$.
@@
This quantity measures kinetic separation between basins, i.e., how difficult it is to transition from one state to another under the Markovian dynamics.
@@
This value is particularly important for identifying deep kinetic traps, which correspond to basins with large self-transition probabilities $T_{jj}$. All temporal quantities (e.g., MFPT, $\tau_{j}$) are scaled by the time step (i.e. $\Delta t=0.1$ ns) to convert frame units into real time in a physical world.
@@ 2.4 Thermodynamic Integrations
While PCA and KTN characterize how post-translational modifications (PTMs) alter the conformational ensemble and kinetics of the protein, they do not directly quantify the energetic cost of introducing a given PTM. To complement these structural and dynamical insights, we perform thermodynamic integration (TI)—a rigorous method from statistical mechanics used to compute the free-energy difference $\Delta G$ between the unmodified reference system and its PTM-perturbed counterpart.
Thermodynamic integration introduces a continuous coupling parameter $\lambda\in[0,1]$ that interpolates between two Hamiltonians, $U(0)$ corresponding to the unmodified (reference) system and $U(1)$ representing the fully perturbed (PTMed) system. For a fixed $\lambda_{k}$, the system is equilibrated and sampled under the hybrid potential $U(\lambda_{k})$, which smoothly blends contributions from both endpoints. The free-energy derivative with respect to $\lambda$, given by the ensemble average of the Hamiltonian gradient,
@@
[…]
@@
Together, TI and BAR provide a rigorous thermodynamic complement to our structural and kinetic analyses, allowing us to assign energetic penalties or stabilizations associated with PTM-induced alterations. These energetic values serve as a quantitative foundation for interpreting how different PTMs influence protein function, binding, or stability.
@@ 3.1 The assembly of the dark complex remains relatively stable after thiol PTMs
We evaluated the global dynamics of the dark complex in its native configuration (Ref 0) and 15 other instances under thiol modifications as described in Table 1. We calculated the temporal evolution of RMSD as well as Rg over 100ns in 3A and 3B, respectively. Ref 1-3 represent the instances of the dark complex with an oxidized thiol on CYS77 with SNO, SSH, or SSG. This CYS77 is equivalent to CYS78 of the GAP2 tetramers in our previous study based on the PDB structure of 2D2I from another cyanobacteria. CYS77 locates near the surface of GAP2 in close proximity to NAD cofactor and CP12. Its oxidation from SNO, SSH, to SSG, signifying a gradual increase in the excluded volume of the side chain. Ref 4-9 represent the reduced thiols with SNO, SSH, or SSG on the cysteine pair, breaking the constraints of disulfide bridges in PRK dimers. Likewise, Ref 10-15 represent the reduced thiols of a disulfide bridge in CP12, relaxing a hairpin confirmation.
While the RMSD profiles vary with PTM instances, their values slowly grow while Rg profiles remain plateaued, indicating that the size of a megacomplex remains steady over time despite the structural perturbation by the inclusion of PTMs at critical sites. The molecular megacomplex maintains its ternary state, rather than undergoing significant conformational changes to a more expanded or collapsed form upon thiol PTMs within 100 ns in the trajectories.
Figure 3: (A) The temporal evolution of backbone RMSD reveals a slight deviation from its starting conformation with 15 kinds of PTM perturbations described in Table 1. Ref 0 is the native state with two built-in disulfide bridges in PRK and two in CP12. A PTM perturbation can be either slightly oxidizing a specific cysteine of GAP2 (Ref 1-3), or breaking disulfide bridges by replacing them with other thiol PTMs in PRK (Ref 4-9) as well as CP12 (Ref 10-15). A plateau suggests reaching equilibrium. (B) Radius of gyration (Rg) over time captures global dynamics reaching a steady state of consistent ternary structures with PTMed residues.

### v2

@@ 15      CYX19, CYX29 SSG
Table 1: Each simulation (Sim.) number corresponds to a specific case with a particular set of PTMs in specific chain locations. “CYS” denotes a free cysteine residue, while “CYX” represents part of a disulfide bridge in PDB ID 6GVE (i.e. Sim.0, and is the considered as the reference or baseline in this work) – there are two disulfide bridges in PRK and in CP12, respectively. Abbreviations of PTMs on a specific cysteine include SNO (S-nitrosylation), SSH (S-sulfhydration), SSG or (S-glutathionylation).
@@ 2.3.1 Univariate Metrics
RMSD quantifies the deviation of instantaneous atomic positions from a reference (average) conformation:
@@
where $\mathbf{r}_{k}^{\rm ref}$ is the position of atom $k$ in the reference state.
@@
RMSF describes per-residue flexibility. For residue $i$, using either backbone heavy-atom center-of-mass or Cα position as $\mathbf{R}_{i}(t)$,
@@
Secondary Structure
Secondary structure assignment tracks helix, sheet, and coil fractions over time using the Dictionary of Secondary Structure in Proteins (DSSP) algorithm26, 27.
@@ 2.3.3 Multivariate Analysis
Multivariate analysis refers to the statistical analysis of data that involves multiple variables or “features”. This type of analysis is essential when dealing with complex and massive datasets where each observation consists of multiple measurements. We deployed multivariate analysis to understand the relationships between the variables in the datasets. We used the technique of principal component analysis (PCA) to reduce the dimensionality of the datasets by transforming them into a new set of orthogonal variables that capture the most variance. By pooling measurements per frame into one feature matrix, we revealed correlated motions invisible to one-dimensional plots by applying PCA, reducing the redundancy among overlapping metrics, and transforming the data from all instances under various PTM perturbations into a common coordinate system28. We ranked the impact of PTM perturbations on the conformational ensemble in the simulations by ranking the variability obtained through PCA.
Feature extraction: For each simulation system $i$ in Table 1, there are $N\approx 1000$ samples. At timeframe $t$, we compute an $M$-dimensional feature vector from a selection of atoms that is either from the whole protein complex or from the reduced representation as described in the previous subsection:
@@
Here, $R_{g}$ and $\mathrm{RMSD}$ are measured in nm. $B$ is the average B-factor measured in nm2, calculated by
@@
where $\langle\mathrm{RMSF}^{2}\rangle$ denotes the average squared RMSF per residue in a selected representation. The $\phi$ and $\psi$ are the average backbone dihedral angles expressed in radians. $\mathrm{SS}_{\mathrm{helix}}$, $\mathrm{SS}_{\mathrm{sheet}}$, and $\mathrm{SS}_{\mathrm{coil}}$ are the fractional secondary-structure propensity. The $\mathrm{n}_{\mathrm{PTM\text{-}Residue}}$ denotes the count of residual contacts between the PTMed residue and its neighboring residues within 3.0 Å.
@@
and concatenating all of 16 simulation instances (e.g., Sim.0 to 15 in Table 1):
@@
and we retain the first $K$ number of principal components such that the cumulative variance preserves, at least, 90% of the total variance,
@@
Ranking PTM perturbations: the mean projection onto the first principal component (PC1) for Sim.ii is given by
$\langle\mathrm{PC1}\rangle_{i}=\frac{1}{N}\sum_{t=1}^{N}(X_{i}\,v_{1})_{t}$ (11)
@@
measures the PTM perturbations on the conformational ensemble from Sim.ii along the dominant mode of variation relative to those of Sim.0. The magnitude $|\Delta\langle\mathrm{PC1}\rangle_{i}|$ indicates deviation from the unmodified reference ensemble of Sim.0, and the sign reflects the direction of the shift in collective conformational space. Simulation systems are ranked by $\Delta\langle\mathrm{PC1}\rangle_{i}$.
The sign of an eigenvector component from $\langle\mathrm{PC1}\rangle_{i}$ informs the importance of individual attributed features: a positive score arises from positive loadings $\ell_{k}$ in $v_{1}$; likewise, negative.
$\langle\mathrm{PC1}\rangle_{i}=\sum_{k=1}^{M}\ell_{k}\,\bar{x}_{i,k}=\mathbf{v}_{1}^{\mathsf{T}}\,\bar{\mathbf{x}}_{i}$ (13)
Here, $v_{1}=[\ell_{1},\ell_{2},\dots,\ell_{M}]^{\mathsf{T}}$ and each loading $\ell_{k}$ informs the strength of the feature $k$ driving first principal component PC1 in either a positive “+” or a negative “–” direction.
2.3.4 Kinetic Transition Network
The kinetic transition network (KTN) formalism transforms a high-dimensional, noisy molecular trajectory into a simplified yet informative network model that captures both the thermodynamics (through stationary distributions) and kinetics (through transition rates and mean first passage time (MFPTs)) of the system. We employed KTN – a graph-based model29– to estimate transition probabilities between metastable states in a conformational ensemble and to identify representative states for each basin. It captures the dynamic connectivity for visualizing and quantifying the kinetics of conformational switching between basins. For this study, we reduced the dimensionality of its feature matrix $X\in\mathbb{R}^{N\times M}$ from Equation (6) by projecting each frame onto the top two principal components. This step transforms the high-dimensional feature space into an interpretable two-dimensional subspace that preserves the most dominant collective motions:
@@ 2.3.4 Kinetic Transition Network
where $\mu_{j}\in\mathbb{R}^{2}$ is the centroid of cluster $j$, and $k$ is a user-defined parameter for defining the number of clusters (typically $k=5$) representing the number of coarse-grained states. Each frame $t$ is assigned a discrete cluster label $m_{t}\in\{0,1,\dots,k-1\}$, identifying the kinetic basin to which it belongs.
@@
In our KTN figures, the node area scales with $\pi_{j}$, and the label “pp” on each node is the stationary probability $\pi_{j}$.
@@
This quantity measures kinetic separation between basins; that is, how difficult it is to transition from one state to another under the Markovian dynamics.
@@
This value is particularly important for identifying deep kinetic traps, which correspond to basins with large self-transition probabilities $T_{jj}$. All temporal quantities (e.g., MFPT, $\tau_{j}$) are scaled by the time step (i.e., $\Delta t=0.1$ ns) to convert frame units into real time in a physical world.
@@ 2.4 Thermodynamic Integrations
While PCA and KTN characterize how PTMs alter the conformational ensemble and kinetics of the protein, they do not directly quantify the energetic cost of introducing a given PTM. To complement these structural and dynamical insights, we perform thermodynamic integration (TI)—a rigorous method from statistical mechanics used to compute the free-energy difference $\Delta G$ between the unmodified reference system and its PTM-perturbed counterpart.
Thermodynamic integration introduces a continuous coupling parameter $\lambda\in[0,1]$ that interpolates between two Hamiltonians, $U(0)$ corresponding to the unmodified reference system and $U(1)$ representing the fully perturbed (PTMed) system. For a fixed $\lambda_{k}$, the system is equilibrated and sampled under the hybrid potential $U(\lambda_{k})$, which smoothly blends contributions from both endpoints. The free-energy derivative with respect to $\lambda$, given by the ensemble average of the Hamiltonian gradient,
@@
$\Delta G=\int_{0}^{1}\left\langle\frac{\partial U}{\partial\lambda}\right\rangle_{\lambda}\,d\lambda.$ (22)
@@
Together, TI and BAR provide a rigorous thermodynamic complement to our structural and kinetic analyses, allowing us to assign energetic penalties or stabilizations associated with PTM-induced alterations. These energetic values serve as a quantitative foundation for interpreting how different PTMs influence protein function, binding, or stability. To calculate relative binding free energy (RBFE), we considered a closed thermodynamic cycle composed of two alchemical transformations:
$\Delta\Delta G_{RBFE}(A\xrightarrow{}B)=\Delta G_{protein}(A\xrightarrow{}B)-\Delta G_{water}(A\xrightarrow{}B)$} (24)
where $\Delta G_{protein}(A\xrightarrow{}B)$ and $\Delta G_{water}(A\xrightarrow{}B)$ represent the transformation from the PTMed cysteine (A) to its regular form (B) in protein complex and bulk water, respectively. For the transformation in bulk water, $\Delta G_{water}(A\xrightarrow{}B)$, we simulated both the cysteine and its glutathionylated form, capped with an acetyl group (N-terminus) and N-methyl group (C-terminus) in TIP3P water. The same simulation protocols used for the protein systems were applied and TI calculations were performed to obtain $\Delta G_{water}(A\xrightarrow{}B)$. For each transformation, we employed the dual topology approach, in which the parameters of the PTMed cysteine in state A were gradually transformed into those of a regular cysteine in state B. Specifically, the sulfur atom from glutathione, covalently bound to the cysteine sulfur, was converted into a hydrogen atom bound to the cysteine sulfur during the alchemical transformation, while the remaining atoms of glutathione were converted into dummy atoms that retained the mass of a hydrogen atom but had no non-bonded interactions. All bonded and non-bonded parameters for state A (PTMed cysteine) were derived using NWChem within the PTM-Psi package.12 This approach has been applied to investigate free energy differences in similar contexts, including point mutations of amino acids in protein complexes,30, 31, 32 pKa estimations of protonated amino acids,33 and ligand binding energies in protein binding pockets.34, 35 Comprehensive details of the TI protocol and parameterization procedures have been reported in our previous study.12
@@ 3 Results and Discussion
3.1 Softwarization: Making PTM-Psi cloud compatible
PTM-Psi is a Python package used to facilitate the computational investigation of PTMs on protein structures and their impacts on dynamics and functions in a workflow that uses well-established software, including AlphaFold2, NWChem, AutoDock, and GROMACS (Figure 2 B). The first version of PTM-Psi assumed that the user was targeting a homogeneous HPC cluster with no GPU accelerators available. To make PTM-Psi a cloud-compatible package, we implemented the “workflow of workflows” approach, wherein a parent workflow triggers one or more child workflows, manages them, and acts on their results. The refactoring into a cloud compatible package results in optimized resource allocation tailored to the computational needs of each child workflow. For example, the most general “workflow of workflows” to predict the free energy gain/penalty associated with the introduction of a PTM in protein involves the following child processes:
  1. 1.
    Protein structure inference from sequence (AlphaFold2)
  2. 2.
    Parametrization of new non-standard amino acid (NWChem)
  3. 3.
    Phase-space sampling through molecular dynamics (MD) simulation (GROMACS)
  4. 4.
    Trajectory analysis and selection of most relevant cases (PTM-Psi analysis module)
  5. 5.
    Thermodynamic integration for alchemical free energy calculations of selected cases (GROMACS)
The first version of PTM-Psi was able to perform steps 1–3 in a semiautomatic fashion using the same number and node types for each step. However, the parametrization of a new non-standard amino acid in NWChem runs only on CPU and has rather large memory demands. In contrast, sampling the phase-space using MD simulations in GROMACS can be performed very efficiently in GPU-resident mode. The new version of PTM-Psi is more automatic and has better control over the resource consumption by spawning child processes that target different hardware architectures. Even better resource optimization could be achieved by letting some of the child workflows mentioned above spawn their own child workflows. For example, the preparation steps needed as prerequisites for the phase-space sampling in GROMACS can be performed in low-cost CPU-only VMs. This type of fine-grained control is an ongoing development.
To take advantage of the heterogeneous nature of cloud-based HPC solutions, we added a (Virtual) machine class to PTM-Psi where the user can specify the number of CPUs, GPUs, and memory available in each VM node. To further enhance user experience in AQE, built-in presets for several Azure VM types were also added to the machine class.
PTM-Psi was also extended with a queue class that can be selected to generate either SLURM or PBSPro batch files that take care of performing all the steps in a workflow and spawning child processes, while tracking the dependencies among different steps.
Finally, we included an additional module to analyze MD trajectories, especially tailored for massive datasets and large macromolecular complexes (see Subsection Workflow of Trajectory Analysis ).
Future enhancements to PTM-Psi capabilities include the addition of a tool that advises users on the best target hardware based on the specific workflow, system size, and resource availability, as well as an interface to the Flux management framework for better job orchestration.
All the capabilities described above are designed to minimize user intervention. However, there are still steps that have not been completely automated. For instance, the workflow to generate new forcefield parameters generates charges, and bond and angle force constants in an automatic fashion, but the user still needs to launch another process to optimize torsional potentials and, subsequently, update the necessary forcefield files manually. Once the forcefield is at hand, the rest of the thermodynamic integration workflow is completely automated, with only one command necessary to queue all necessary child workflows.
3.2 Hardware Considerations and the Need for PTM-Psi on the Cloud
PTM-Psi relies heavily on GROMACS9 and NWChem8 for the most computationally demanding tasks. Both software stacks have different capabilities in terms of the hardware requirements and scalability of the algorithms being used. As a consequence, adaptation of each part of the workflows to the appropriate hardware is needed.
The choice of computational platform involves trade-offs between capability, accessibility, and timeliness. In traditional on-premise HPC clusters, homogeneity was the dominant design philosophy. This led to systems with a large number of nodes with (nearly) identical hardware. Although this model has some advantages, like simpler management and more predictable performance, it can also lead to a lack of flexibility in the software stack and/or to non-optimal performance. Moreover, deploying and expanding an on-premise HPC cluster is a complex, multi-staged project with a rather long (months to years) timeframe to complete, leading to slow procurement cycles, limited hardware diversity, and usually long queue times.
In contrast, HPC cloud computing is the on-demand delivery of computing power and other information technology resources via the internet with a pay-as-you-go pricing scheme. Large cloud service providers (CSP) like Microsoft Azure, Amazon Web Services, and Google Cloud Platform offer virtually limitless computing power distributed across their multi-geographic region data centers. They democratize access to a vast and varied landscape of cutting-edge hardware, bypassing the formidable entry barriers to Leadership Computing Facilities and the slowness of on-premise HPC procurement. Both HPC, in which a large number of processors are needed to solve one particularly complex task, and high-throughput computing (HTC), in which a large number of simple and computationally independent tasks are needed, can be tackled with HPC cloud services. Since CSPs usually offer a large set of different resources, the hardware can be configured to perfectly suit the application requirements, allowing for the perfect match between hardware and workload. Crucially, the on-demand nature of the cloud services eliminates queueing delays–a decisive advantage for our work.
The elasticity and heterogeneity of Cloud resources are ideal for the workflows implemented in PTM-Psi. For instance, while AlphaFold2 and GROMACS can be run using NVIDIA GPUs, only GROMACS officially supports AMD and Intel GPUs. We also note that non-standard residue parametrization with NWChem can only run on CPUs. For these reasons, making PTM-Psi cloud-compatible, by leveraging hardware heterogeneity and exploiting the “workflow of workflows” approach, might decrease turnaround times and could potentially lead to cost savings.
As a showcase of the efficiencies that can be obtained with PTM-Psi on the cloud, we mention that we were able to run a full workflow of thermodynamic integration for 200 PTM combinations in the dark complex (800k atoms) in less than 72 hours of active cloud allocation. This time includes the generation of forcefield parameters with NWChem using 20 HBv2 VMs, phase-space sampling with GROMACS using 20 NC_H100 VMs, and the trajectory analysis and thermodynamic integration using 20 HBv2 VMs without human supervision between each child step. The NC_H100 VMs, with a higher cost, were allocated only for 48 hours. Based on historical utilization from other on-premise clusters, we estimate that the same simulation campaign would have taken several weeks to complete.
Trajectory analysis and thermodynamic integration results for a relevant subset of the 200 combinations are presented in the following subsections.
@@ 3.3 The assembly of the dark complex remains relatively stable after thiol PTMs
We evaluated the global dynamics of the dark complex in its native configuration (Sim.0) and 15 other instances under thiol modifications as described in Table 1. We calculated the temporal evolution of RMSD as well as $R_{g}$ over 100ns in Figure 3A and 3B, respectively. Sim.1-3 represent the instances of the dark complex with an oxidized thiol on CYS77 with SNO, SSH, or SSG. This CYS77 is equivalent to CYS78 of the GAP2 tetramers in our previous study based on the PDB structure of “2D2I” from another cyanobacteria. CYS77 locates near the surface of GAP2 in close proximity to NAD cofactor and CP12. Its oxidation from SNO and SSH, to SSG, signifying a gradual increase in the excluded volume of the side chain. Sim.4-9 represent the reduced thiols with SNO, SSH, or SSG on the cysteine pair, breaking the constraints of disulfide bridges in PRK dimers. Likewise, Sim.10-15 represent the reduced thiols of a disulfide bridge in CP12, relaxing a hairpin conformation.
While the RMSD profiles vary with PTM instances, their values slowly grow while $R_{g}$ profiles remain plateaued, indicating that the size of a megacomplex remains steady over time despite the structural perturbation by the inclusion of PTMs at critical sites. The molecular megacomplex maintains its ternary state, rather than undergoing significant conformational changes to a more expanded or collapsed form upon thiol PTMs within 100 ns in the trajectories.
Figure 3: (A) The temporal evolution of backbone RMSD reveals a slight deviation from its starting conformation with 15 kinds of PTM perturbations described in Table 1. Sim.0 is the native state with two built-in disulfide bridges in PRK and two in CP12. A PTM perturbation can be either slightly oxidizing a specific cysteine of GAP2 (Sim.1-3), or breaking disulfide bridges by replacing them with other thiol PTMs in PRK (Sim.4-9) as well as CP12 (Sim.10-15). The plateau reflects that conformational sampling has converged toward equilibrium during the MD simulations. (B) The $R_{g}$ over time captures global dynamics reaching a steady state of consistent ternary structures with PTMed residues.